OpenStudy (hellokitty17):

Simplify completely the quantity 8 times x to the third power plus 6 times x to the 2nd power minus 4 times x all over negative 2 times x. a.-4x2 - 3x + 2 b.-4x2 + 3x + 2 c.4x4 + 3x3 - 2x2 d.4x2 + 6x2 - 4x

OpenStudy (tkhunny):

\(\dfrac{a+b+c}{d} = \dfrac{a}{d}+\dfrac{b}{d}+\dfrac{c}{d}\)

oregonduck (oregonduck):

you make no sense that is confusing

OpenStudy (hellokitty17):

I think it is a

oregonduck (oregonduck):

yup you are right 8x^3 +6x^2 -4x ---------------- = - 4x^2 - 3x +2 - 2x so is a . right

OpenStudy (tkhunny):

And just how did you do that, @OregonDuck ?

oregonduck (oregonduck):

i did it because i am smart

OpenStudy (tkhunny):

You did not answer my question. I didn't say "Why did you do it?" or "Wow, that's amazing. You must be the smartest person on Earth." I just said, "How did you do that?" Where, exactly, did "-4x^2" come from, for example?